# Gaussian curvature and Lyapunov exponent as probes of black hole phase transitions

### v1

@@ Abstract
We study the Gaussian curvature of unstable null orbits. The Gaussian curvature exhibits multivaluedness near the phase transition point of a first-order phase transition. Numerical investigations of Reissner-Nordström Anti-de Sitter (RN-AdS), Hayward-AdS, and Hayward-Letelier-AdS black holes demonstrate that this geometric multivalued region coincides precisely with the spinodal region calculated by black hole thermodynamics. Using the known relation $K=-\lambda^{2}$ linking orbital geometry to chaotic dynamics, we show that this geometric feature also satisfies the critical exponents predicted by mean-field theory, consistent with those derived from Lyapunov exponents. Our work demonstrates that Gaussian curvature can serve as an alternative effective tool to study the phase structure of black holes.
@@ I Introduction
In black hole thermodynamics, the multivaluedness of QNMs Liu et al. (2014); Mahapatra (2016); Chabab et al. (2016); Zou et al. (2017); Zhang et al. (2021); Zhao et al. (2023, 2024, 2025a, 2025b); Hou et al. (2025) and Lyapunov exponents Guo et al. (2022b); Yang et al. (2023); Lyu et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); Karthik R. and Hegde (2025); Awal and Phukon (2025); Yang et al. (2025a) during first-order phase transitions suggest a potential connection between these quantities and black hole phase transitions. Studies of the observable photon sphere have also tied it to black hole phase transitions Wei and Liu (2018); Wei et al. (2019); Zhang et al. (2019); Xu et al. (2019); Li et al. (2020a); Naveena Kumara et al. (2020); Du et al. (2023); Kumar et al. (2024a); Yang et al. (2025b), confirming that dynamical analysis can probe such phenomena. However, one question remains largely unexplored. Since general relativity is fundamentally geometric, how does spacetime geometry itself change during a thermal phase transition? Is there an intrinsic geometric quantity that can directly reflect this transition?
Black holes are exact solutions of Einstein’s equations whose physical processes are reflected by changes in geometric properties. Gaussian curvature, an intrinsic measure of a two-dimensional manifold, directly quantifies local spacetime curvature and serves as an ideal probe of geometry. Using Gaussian curvature, researchers have developed purely geometric methods to study black hole photon spheres Qiao and Li (2022); Chern (2006), thus establishing a clear correspondence between dynamics and geometry. Recently, it has been used to analyze the stability of circular orbits for massless particles, to determine the existence and distribution of stable and unstable orbits Qiao (2025), and to investigate gravitational lensing by massive particles Li et al. (2020b); Qiao and Zhou (2023). A key finding is that the Lyapunov exponent of null circular orbits near black holes is linked to their Gaussian curvature Gallo and Mädler (2025). This raises a previously overlooked question: can geometric quantities such as Gaussian curvature be used to connect with first-order phase transitions of a black hole?
@@
This paper is arranged as follows. In Sec. II we briefly review the relations of black hole thermodynamics, the Gaussian curvature of two-dimensional surfaces, and the calculation of Lyapunov exponents. In Sec. III we present the investigation of three distinct black holes and the analysis of the results. In Sec. IV we address the computation of critical exponents. In Sec. V we offer a summary and discussion, and the appendix follows. We set $G=c=k_{B}=\hbar=1$ in this paper.
@@ II Theoretical Framework
In this section we first briefly review the relations of black hole thermodynamics, then introduce how to obtain the Gaussian curvature from a two-dimensional Riemannian metric, and finally recall the derivation of the Lyapunov exponent and its connection to the Gaussian curvature.
@@ II.1 Thermodynamics and phase structure of spherically symmetric black holes
In this subsection we briefly review the relations in black hole thermodynamics by considering a 3+1-dimensional spherically symmetric black hole solution
@@
Note that $f(r_{+})=0$, where $r_{+}$ is the horizon radius. The Hawking temperature can be obtained directly from the metric as
$T=\frac{f^{\prime}(r_{+})}{4\pi}.$ (2)
The Gibbs free energy is
@@ II.2 Gaussian curvature of two-dimensional manifolds
Gaussian curvature $K$ is an intrinsic geometric quantity of the optical metric associated with massless particles, and the light ring is closely linked to black hole spacetime. This suggests that $K$ on the light ring may exhibit anomalous behavior near the phase transition point, offering a potential geometric signature for studying such transitions.
Consider a general 3+1-dimensional spherically symmetric metric
@@
For the unstable null circular orbit ($r=r_{LR}$, see the appendix for details), we have the following formula
$f^{\prime}(r_{LR})=\frac{2f(r_{LR})}{r_{LR}}.$ (11)
Inserting Eq. (11) into (10) yields a known relation Gallo and Mädler (2025)
@@
In Sec. III we will demonstrate the connection between Gaussian curvature and first-order phase transitions by examining the anomalous behavior of $K$ for circular null orbits near three distinct black holes. A more explicit link between the detailed behavior of $K$ and the phase transition will be provided in the next subsection through the analysis of Lyapunov exponents.
@@ II.3 Lyapunov exponent for unstable orbits
where $\sigma=\{+1,-1,0\}$ denotes timelike, spacelike, and null geodesics, respectively. The angular momentum and energy are defined by
@@
$(\dot{\varepsilon})^{2}-\frac{1}{2}V^{\prime\prime}_{\text{eff}}(r_{0})\varepsilon^{2}=0.$ (18)
Note that this is an inverted harmonic oscillator equation, defining the Lyapunov exponent Bhattacharyya et al. (2021); Cardoso et al. (2009)
$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V^{\prime\prime}_{\text{eff}}(r_{0})}.$ (19)
@@
Eq. (21) establishes a quantitative relation between Gaussian curvature and Lyapunov exponent on the light ring. Previous studies Guo et al. (2022b); Yang et al. (2023); Lyu et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); Karthik R. and Hegde (2025); Awal and Phukon (2025) have demonstrated that $\lambda$ becomes multivalued near the phase transition point. In Refs.  Guo et al. (2022b); Shukla et al. (2024) the authors interpreted the Lyapunov exponent as an order parameter that decreases with increasing temperature $T$. Clearly, Eq. (21) provides a mapping between $K$ and $\lambda$ on the light ring. Such a mapping paves the way for a connection between geometry and thermodynamics, thereby enabling the investigation of thermodynamics from a geometric perspective. Combined with Eq. (21), if $\lambda$ is multivalued in the spinodal region, $K$ must also be multivalued, offering a geometric way to probe black hole phase transitions. Note that not only do null circular orbits admit a Lyapunov exponent, but circular orbits of massive particles can also be defined. In both cases, these dynamical quantities connect with the phase transition.
For completeness, we also examine the Lyapunov exponent associated with the circular orbits of massive particles. The corresponding unstable circular orbit is located at $r=r_{0}$ (see Appendix) with $\sigma=1$, and its Lyapunov exponent is given by
@@ III Numerical Results and Analysis
III.1 RN-AdS black hole
We begin with the RN-AdS black hole. It is a static, charged, four-dimensional spherically symmetric solution whose metric function is
$f_{R}(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{\ell^{2}},$ (24)
where $Q$ denotes the black hole’s charge, and $\ell$ is the AdS radius. The Hawking temperature is
$T_{R}=\frac{1}{4\pi r_{+}}\left(1-\frac{Q^{2}}{r_{+}^{2}}+\frac{3r_{+}^{2}}{\ell^{2}}\right).$ (25)
Its Gibbs free energy is given by
$F_{R}=\frac{1}{4}\left(\frac{3Q^{2}}{r_{+}}+r_{+}-\frac{r_{+}^{3}}{\ell^{2}}\right).$ (26)
Here, we use the relation
$M=\frac{r_{+}}{2}\left(1+\frac{Q^{2}}{r_{+}^{2}}+\frac{r_{+}^{2}}{\ell^{2}}\right).$ (27)
By dimensional analysis, we introduce the following scaling
$\tilde{r}_{+}=\frac{r_{+}}{\ell},\quad\tilde{Q}=\frac{Q}{\ell},\quad\tilde{M}=\frac{M}{\ell},\quad\tilde{F}_{R}=\frac{F_{R}}{\ell},\quad\tilde{T}_{R}=T_{R}\ell.$ (28)
The multivalued $\lambda_{R}(\tilde{T}_{R})$ near the phase transition point of RN-AdS black holes is well-established Kubiznak and Mann (2012); Guo et al. (2022b). Here, we present only the behavior of the Gaussian curvature $K_{R}$ for its light rings at $\tilde{T}_{p1}$.
Figure 1: (a) Free energy $\tilde{F}_{R}$ and (b) Gaussian curvature $K_{R}$ of unstable photon orbits versus temperature $\tilde{T}_{R}$ for RN-AdS black holes, and $\tilde{Q}=\frac{1}{8.66},~\tilde{Q}_{c}=\frac{1}{6},~\tilde{Q}<\tilde{Q}_{c}$. Swallowtail structures in $\tilde{F}_{R}$ and multivalued $K_{R}$ at $\tilde{T}_{p1}$ evidence geometric degeneracy during phase transitions.
As shown in Fig. 1, in the canonical ensemble (fixed charge $\tilde{Q}$) and when $\tilde{Q}<\tilde{Q}_{c}$, $\tilde{F}_{R}(\tilde{T}_{R})$ shows the characteristic swallowtail structure, which means a first-order phase transition appears, whose transition point is located at $\tilde{T}_{p1}$. Consequently, in Fig. 1, within the spinodal region $\tilde{T}_{R}\in(\tilde{T}_{R1},\,\tilde{T}_{R2})$, the curve $K_{R}(\tilde{T}_{R})$ exhibits multivalued behavior. This occurs because three black hole solutions (small black hole, intermediate black hole, and large black hole) exist in this region. Consequently, a fixed temperature $\tilde{T}_{R}$ corresponds to multiple values of $K_{R}$, demonstrating that the Gaussian curvature $K_{R}$ is multivalued in the spinodal region.
III.2 Hayward–AdS black hole
Unlike the RN-AdS black hole, the Hayward-AdS black hole is a regular solution in AdS spacetime that avoids the appearance of a spacetime singularity J. (1968); Hayward (2006). Its thermodynamics and phase transitions have been investigated in Ref. Gogoi et al. (2024). The metric function is given by
$f_{H}(r)=1-\frac{2Mr^{2}}{g^{3}+r^{3}}+\frac{r^{2}}{\ell^{2}},$ (29)
where $g$ denotes the magnetic monopole charge arising from the nonlinear electrodynamics term in the action Kumar et al. (2024b), and $\ell$ is the AdS radius. The Hawking temperature is
$T_{H}=\frac{\ell^{2}(r_{+}^{3}-2g^{3})+3r_{+}^{5}}{4\pi\ell^{2}r_{+}(g^{3}+r_{+}^{3})}.$ (30)
The corresponding Gibbs free energy reads
$F_{H}=\frac{2(g^{3}+r_{+}^{3})^{2}\left(1+\frac{r_{+}^{2}}{\ell^{2}}\right)-r_{+}^{3}\left[(r_{+}^{3}-2g^{3})+\frac{3r_{+}^{5}}{\ell^{2}}\right]}{4r_{+}^{2}(g^{3}+r_{+}^{3})}.$ (31)
Here, we use the relation
$M=\left(1+\frac{r_{+}^{2}}{\ell^{2}}\right)\frac{g^{3}+r_{+}^{3}}{2r_{+}^{2}}.$ (32)
We introduce the following scaling
Figure 2: (a) Free energy $\tilde{F}_{H}$ and (b) Gaussian curvature $K_{H}$ of unstable photon orbits versus temperature $\tilde{T}_{H}$ for Hayward-AdS black holes, and $\tilde{g}=0.061538,~\tilde{g}_{c}=0.142336,~\tilde{g}<\tilde{g}_{c}$. Swallowtail structures in $\tilde{F}_{H}$ and multivalued $K_{H}$ at $\tilde{T}_{p2}$ evidence geometric degeneracy during phase transitions.
Previous work has extensively investigated the thermodynamic properties of the Hayward-AdS black hole, including the multivalued behavior of the Lyapunov exponent $\lambda_{H}(\tilde{T}_{H})$ near $\tilde{T}_{p2}$ Gogoi et al. (2024). Here, we focus on presenting the $K_{H}(\tilde{T}_{H})$ curve for null geodesics and its thermodynamic counterpart, the free energy $\tilde{F}_{H}(\tilde{T}_{H})$. Following the same procedure as in the RN-AdS case, the numerical results are shown in Figs. 2.
Figs. 2 and 2 show that, for the Hayward–AdS black hole, when $\tilde{g}<\tilde{g}_{c}$, the Gaussian curvature $K_{H}$ for null geodesics also exhibits multivalued behavior around the first-order phase transition point, as in the RN–AdS case.
III.3 Hayward-Letelier-AdS black hole
@@
where $g$ is the magnetic monopole charge, $\ell$ is the AdS radius, and the string-cloud parameter $a$ arises from the string-cloud term in action Kumar et al. (2024b). When $a=0$, the metric reduces to the Hayward–AdS black hole. The Hawking temperature is
$T_{HL}=\frac{\left[\ell^{2}(r_{+}^{3}-2g^{3})+3r_{+}^{5}+a\ell^{2}(2g^{3}-r_{+}^{3})\right]}{4\pi\ell^{2}r_{+}(r_{+}^{3}+g^{3})}.$ (35)
@@
$\begin{split}F_{HL}=&\frac{2(g^{3}+r_{+}^{3})^{2}\left(1-a+\frac{r_{+}^{2}}{\ell^{2}}\right)}{4r_{+}^{2}(g^{3}+r_{+}^{3})}\\ &-\frac{r_{+}^{3}\left[(1-a)(r_{+}^{3}-2g^{3})+\frac{3r_{+}^{5}}{\ell^{2}}\right]}{4r_{+}^{2}(g^{3}+r_{+}^{3})}.\end{split}$ (36)
@@
Since $a$ is a dimensionless constant, we use the same scaling Eq. (33). Following the same procedure as before, we present the $\tilde{F}_{HL}(\tilde{T}_{HL})$, $K_{HL}(\tilde{T}_{HL})$, and $\lambda_{HL}(\tilde{T}_{HL})$ curves for the Hayward–Letelier–AdS black hole.
@@
Figure 3: (a) $\tilde{F}_{HL}$, (b) $K_{HL}$ of unstable null orbits and (c-d) $\lambda_{HL}$ of unstable null/timelike orbits versus temperature $\tilde{T}_{HL}$ for Hayward-Letelier-AdS black holes, and $\tilde{g}=0.061538,~\tilde{g}_{c}=0.09002,~a=0.6,~\tilde{g}<\tilde{g}_{c}$. Swallowtail structures in $\tilde{F}_{HL}$ and multivalued $K_{HL}$ at $\tilde{T}_{p3}$ evidence geometric degeneracy during phase transitions. $\lambda_{HL}$ at $\tilde{T}_{p3}$ also exhibits such multivaluedness.
As illustrated in Fig. 3(a), for $\tilde{g}<\tilde{g}_{c}$ the swallowtail structure indicates a first-order phase transition. Similarly, in Figs. 3(b), 3(c), and 3(d), the $K_{HL}(\tilde{T}_{HL})$ curve for null circular orbits and the $\lambda_{HL}(\tilde{T}_{HL})$ curve for both timelike and null orbits exhibit multivalued behavior. This shows that both the Gaussian curvature $K_{HL}$ and the Lyapunov exponent $\lambda_{HL}$ are intimately related to the first-order phase transition, because Eq. (21) connects chaotic dynamics and geometry. This connection paves the way for studying thermodynamics from a geometric perspective.
Moreover, we find that the Gaussian curvature $K_{HL}$ of the unstable null circular orbits in Figs. 12, and 3(b) is consistently negative, in full agreement with the result of Ref. Qiao and Li (2022). In Ref. Qiao and Li (2022), the authors adopted an alternative approach based on the Hadamard theorem, demonstrating that $K<0$ corresponds to unstable circular orbits, while $K>0$ indicates stable ones. Our results show that the Gaussian curvature on the light ring is strictly negative, supporting their conclusion. Additionally, after the black hole undergoes a first-order phase transition, specifically on the right side of the spinodal region, $|K|$ decreases monotonically with increasing temperature $\tilde{T}$. This behavior suggests that $|K|$ may be related to the order parameter, a finding consistent with other recent studies Refs. Guo et al. (2022b); Shukla et al. (2024); Banerjee and Roychowdhury (2012a, b). These results indicate that the Gaussian curvature becomes multivalued whenever the black hole experiences a phase transition.
Figure 4: (a) Free energy $\tilde{F}_{HL}$ and (b) Gaussian curvature $K_{HL}$ versus temperature $\tilde{T}_{HL}$ for a Hayward-Letelier-AdS black hole without phase transition ($\tilde{g}=0.092308,~\tilde{g}_{c}=0.09002,~a=0.6,~\tilde{g}>\tilde{g}_{c}$). Both quantities exhibit monotonic behavior, confirming the absence of thermodynamic criticality.

### v2

@@ Abstract
First-order phase transitions of black holes have been extensively studied within thermodynamic frameworks, yet the corresponding evolution of spacetime geometric properties remains unclear. This paper establishes a purely differential geometric framework to probe such phase transitions by analyzing the curvature of unstable null orbits. Using the geodesic curvature of the optical metric to locate the light ring, we demonstrate that the corresponding Gaussian curvature $K$ serves as a direct geometric signature of the phase transition. During a first-order phase transition, the curve $K$ versus temperature $T$ exhibits a multivalued structure within the spinodal region, precisely mirroring the swallowtail behavior of the free energy. Numerical analysis of Hayward-Letelier-AdS black holes confirms the effectiveness of this geometric signature. Our work demonstrates that the intrinsic geometric quantities of spacetime encode the information of black hole phase transitions. These quantities serve simultaneously as geometric probes and order parameters for black hole phase transitions. This result provides a purely geometric foundation for understanding the correspondence between thermodynamics and spacetime curvature in the null case.
@@ I Introduction
In black hole thermodynamics, the multivaluedness of QNMs Liu et al. (2014); Mahapatra (2016); Chabab et al. (2016); Zou et al. (2017); Zhang et al. (2021); Zhao et al. (2023, 2024, 2025a, 2025b); Hou et al. (2025) and Lyapunov exponents Guo et al. (2022b); Yang et al. (2023); Lyu et al. (2024); Kumara et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); Karthik R. and Hegde (2025); Awal and Phukon (2025); Yang et al. (2025a) during first-order phase transitions suggest a potential connection between these quantities and black hole phase transitions. Studies of the observable photon sphere have also tied it to black hole phase transitions Wei and Liu (2018); Wei et al. (2019); Zhang et al. (2019); Xu et al. (2019); Li et al. (2020a); Naveena Kumara et al. (2020); Du et al. (2023); Kumar et al. (2024a); Yang et al. (2025b), confirming that dynamical analysis can probe such phenomena. However, one question remains largely unexplored. Since general relativity is fundamentally geometric, how does spacetime geometry itself change during a thermal phase transition? Is there an intrinsic geometric quantity that can directly reflect this phase transition?
Black holes are exact solutions of the Einstein equations whose physical processes are reflected by changes in geometric properties. Gaussian curvature, an intrinsic measure of a two-dimensional manifold, directly quantifies local spacetime curvature and serves as an ideal probe of geometry. Using Gaussian curvature, researchers have developed purely geometric methods to study black hole photon spheres Qiao and Li (2022); Chern (2006), thus establishing a clear correspondence between dynamics and geometry. Recently, it has been used to analyze the stability of circular orbits for massless particles, to determine the existence and distribution of stable and unstable orbits Qiao (2025), and to investigate gravitational lensing by massive particles Li et al. (2020b); Qiao and Zhou (2023). This raises a previously overlooked question: can geometric quantities such as Gaussian curvature be used to connect with first-order phase transitions of a black hole? A key finding is that the Lyapunov exponent of null circular orbits near black holes is linked to their Gaussian curvature Gallo and Mädler (2025). The established role of the Lyapunov exponent as an effective probe makes it possible to address the above question.
@@
This paper is arranged as follows. In Sec. II, we briefly review the relations of black hole thermodynamics, the Gaussian curvature of two-dimensional surfaces, and the calculation of Lyapunov exponents. In Sect. III, we present the investigation of the Hayward-Letelier-AdS black hole and the analysis of the results. In Sec. IV, we address the computation of critical exponents. In Sec. V, we offer a summary and discussion, and the appendix follows. We set $G=c=k_{B}=\hbar=1$ in this paper.
@@ II Theoretical Framework
In this section, we first briefly review the relations of black hole thermodynamics, then introduce how to obtain the Gaussian curvature from a two-dimensional Riemannian metric, and finally recall the derivation of the Lyapunov exponent and its connection to the Gaussian curvature.
@@ II.1 Thermodynamics and phase structure of spherically symmetric black holes
In this subsection, we briefly review the relations in black hole thermodynamics by considering a 3+1-dimensional spherically symmetric black hole solution
@@
Note that $f(r_{+})=0$, where $r_{+}$ is the horizon radius. The temperature can be obtained directly from the first law of a charged and static black hole in extended phase space as Kumar et al. (2024b)
$T=\frac{\partial M}{\partial S},$ (2)
for a fixed charge and pressure. The Gibbs free energy is
@@ II.2 Gaussian curvature of two-dimensional manifolds
Gaussian curvature $K$ is an intrinsic geometric quantity of the optical metric associated with massless particles, and the light ring is closely linked to black hole spacetime. This suggests that $K$ on the light ring may exhibit anomalous behavior near the phase transition point, offering a potential geometric signature for studying such phase transitions.
Consider a static 3+1-dimensional spherically symmetric metric
@@
Consider the optical metric given in Eq. (7), its geodesic curvature $\kappa_{g}$ is given by
$\kappa_{g}=\sqrt{\frac{g(r)}{f(r)}}\frac{2f(r)-rf^{\prime}(r)}{2r}.$ (11)
On the unstable null circular orbit $r=r_{LR}$ (the light ring), the geodesic curvature vanishes, yielding
$2f(r_{LR})=r_{LR}f^{\prime}(r_{LR}).$ (12)
For the unstable null circular orbit ($r=r_{LR}$, inserting Eq. (12) into (10) yields a known relation Gallo and Mädler (2025)
@@
In Sec. III, we will demonstrate the connection between Gaussian curvature and first-order phase transitions by examining the anomalous behavior of $K$ for circular null orbits near the Hayward-Letelier-AdS. A more explicit link between the detailed behavior of $K$ and the phase transition will be provided in the next subsection through the analysis of Lyapunov exponents.
@@ II.3 Lyapunov exponent for unstable orbits
where $\sigma=\{+1,-1,0\}$ denotes spacelike, timelike, and null geodesics, respectively. The angular momentum and energy are defined by
@@
$(\dot{\varepsilon})^{2}-\frac{1}{2}V^{\prime\prime}_{\text{eff}}(r_{LR})\varepsilon^{2}=0,$ (19)
where $\varepsilon=r-r_{LR}$. Note that this is an inverted harmonic oscillator equation, defining the Lyapunov exponent Bhattacharyya et al. (2021); Cardoso et al. (2009)
$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V^{\prime\prime}_{\text{eff}}(r_{LR})}.$ (20)
@@
Eq. (22) establishes a quantitative relation between Gaussian curvature and Lyapunov exponent on the light ring. Previous studies Guo et al. (2022b); Yang et al. (2023); Lyu et al. (2024); Du et al. (2025); Shukla et al. (2024); Gogoi et al. (2024); Chen et al. (2025); Karthik R. and Hegde (2025); Awal and Phukon (2025) have demonstrated that $\lambda$ becomes multivalued near the phase transition point. In Refs. Guo et al. (2022b); Shukla et al. (2024), the authors interpreted the Lyapunov exponent as an order parameter that decreases with increasing temperature $T$. Clearly, Eq. (22) provides a mapping between $K$ and $\lambda$ on the light ring. Such a mapping paves the way for a connection between geometry and thermodynamics, thereby enabling the investigation of thermodynamics from a geometric perspective. Combined with Eq. (22), if $\lambda$ is multivalued in the spinodal region, $K$ must also be multivalued, offering a geometric way to probe black hole phase transitions. Note that not only do null circular orbits admit a Lyapunov exponent, but circular orbits of massive particles can also be defined. In both cases, these dynamical quantities connect with the phase transition.
For completeness, we also examine the Lyapunov exponent associated with the circular orbits of massive particles. The corresponding unstable circular orbit is located at $r=r_{0}$ with $\sigma=-1$, and its Lyapunov exponent is given by
@@ III Numerical Results and Analysis
where $g$ is the magnetic monopole charge, $\ell$ is the AdS radius, and the string-cloud parameter $a$ arises from the string-cloud term in action Kumar et al. (2024b). For studies of Hayward-AdS black holes with string fluids in the extended thermodynamic phase space, see Ref. Anand et al. (2025) for thermodynamic curvature and topological properties, and Ref. Singh et al. (2025) for chaotic behavior. When $a=0$, the metric reduces to the Hayward–AdS black hole. The temperature is Anand et al. (2025)
$T_{HL}=\frac{r_{+}^{3}\left[(1-a)+3r_{+}^{2}/\ell^{2}\right]-2g^{3}(1-a)}{4\pi r_{+}^{4}}.$ (26)
@@
$\begin{split}F_{HL}=\frac{1}{4}\left\{g^{3}\left[\frac{2}{\ell^{2}}+\frac{4(1-a)}{r_{+}^{2}}\right]+(1-a)r_{+}-\frac{r_{+}^{3}}{\ell^{2}}\right\}.\end{split}$ (27)
@@
Since $a$ is a dimensionless constant, We introduce the following scaling
We present the $\tilde{F}_{HL}(\tilde{T}_{HL})$, $K_{HL}(\tilde{T}_{HL})$, and $\lambda_{HL}(\tilde{T}_{HL})$ curves for the Hayward–Letelier–AdS black hole.
@@
Figure 1: Thermodynamic, geometric and chaos signatures of the first-order phase transition in Hayward-Letelier-AdS space-times. (a) $\tilde{F}_{HL}$, (b) $K_{HL}$ of unstable null orbits and (c-d) $\lambda_{HL}$ of unstable null/timelike orbits versus temperature $\tilde{T}_{HL}$ for Hayward-Letelier-AdS black holes, and $\tilde{g}=0.0615,~\tilde{g}_{c}=0.1042,~a=0.6,~\tilde{g}<\tilde{g}_{c},~L=20\ell$. Swallowtail structures in $\tilde{F}_{HL}$ and multivalued $K_{HL}$ at $\tilde{T}_{p}$ evidence geometric degeneracy during phase transitions. $\lambda_{HL}$ at $\tilde{T}_{p}$ also exhibits such multivaluedness.
As illustrated in Fig. 1(a), for $\tilde{g}<\tilde{g}_{c}$ the swallowtail structure indicates a first-order phase transition. Similarly, in Figs. 1(b), 1(c), and 1(d), the $K_{HL}(\tilde{T}_{HL})$ curve for null circular orbits and the $\lambda_{HL}(\tilde{T}_{HL})$ curve for both timelike and null orbits exhibit multivalued behavior. This shows that both the Gaussian curvature $K_{HL}$ and the Lyapunov exponent $\lambda_{HL}$ are intimately related to the first-order phase transition, because Eq. (22) connects chaotic dynamics and geometry. This connection paves the way for studying thermodynamics from a geometric perspective.
Moreover, we find that the Gaussian curvature $K_{HL}$ of the unstable null circular orbits in Fig. 1(b) is consistently negative, in full agreement with the result of Ref. Qiao and Li (2022). In Ref. Qiao and Li (2022), the authors adopted an alternative approach based on the Hadamard theorem, demonstrating that $K<0$ corresponds to unstable circular orbits, while $K>0$ indicates stable ones. Our results show that the Gaussian curvature on the light ring is strictly negative, supporting their conclusion. Additionally, after the black hole undergoes a first-order phase transition, specifically on the right side of the spinodal region, $|K|$ decreases monotonically with increasing temperature $\tilde{T}$. This behavior suggests that $K$ may be related to the order parameter, a finding consistent with other recent studies Refs. Guo et al. (2022b); Shukla et al. (2024); Banerjee and Roychowdhury (2012a, b). These results indicate that the Gaussian curvature becomes multivalued whenever the black hole experiences a phase transition.
Figure 2: Thermodynamic and geometric signatures of the absence of a first-order phase transition in Hayward-Letelier-AdS space-times. (a) Free energy $\tilde{F}_{HL}$ and (b) Gaussian curvature $K_{HL}$ versus temperature $\tilde{T}_{HL}$ for a Hayward-Letelier-AdS black hole without phase transition ($\tilde{g}=0.0615,~\tilde{g}_{c}=0.1042,~a=0.6,~\tilde{g}>\tilde{g}_{c}$). Both quantities exhibit monotonic behavior, confirming the absence of thermodynamic criticality.